# Safe and Efficient Social Navigation through Explainable Safety Regions Based on Topological Features

### v1

@@ 2 Related work
Moreover, persistent entropy has demonstrated its effectiveness in a variety of other applications such as characterizing idiotypic immune networks [25], analyzing similarities in piecewise linear functions and time series [7], and separating topological features from noise in Vietoris-Rips complexes [26]. These successful applications underscore its versatility and potential to provide robust insights across diverse domains.
@@ 3.1 Topology Background
Topology and Topological Space. A topological space is a powerful mathematical concept for describing the connectivity of a space. Informally, a topological space is a set of points, each of them equipped with the notion of neighbourhood.
@@
Simplicial Complex. An abstract simplicial complex $\mathcal{K}$ is given by:
  • A set $V$ of 00-simplices (also called vertices);
@@
A simplicial complex $K$ is a geometric realization of an abstract simplicial complex $\mathcal{K}$ [29]. It is constructed as a nested family of simplices, where each simplex is a generalization of geometric shapes of various dimensions. Specifically, a 00-simplex corresponds to a point in a given Euclidean space $\mathbb{R}^{n}$, a $1$-simplex represents an edge, a $2$-simplex is a filled triangle, a $3$-simplex is a filled tetrahedron, and so on. This hierarchical structure allows the simplicial complex to model higher-dimensional relationships between points in a mathematically rigorous way.
[…]
@@
Persistent Homology. Topological data analysis, particularly persistent homology, has emerged as a valuable approach for studying the geometry and connectivity of datasets that evolve over time or space [27]. The primary objective of employing topological data analysis tools, like persistent homology and persistent entropy, in the context of fleet behavior modeling is to achieve a more nuanced understanding of system dynamics.
@@
Persistent homology describes how the homology of $K$ changes along a filtration $\{K(t)\mid t\in\mathbb{R}\}$. A $k$-dimensional Betti interval, with endpoints $[t_{\text{start}},t_{\text{end}})$, corresponds to a $k$-dimensional hole that appears at filtration time $t_{\text{start}}$ and remains until time $t_{\text{end}}$.
To visualize the persistence of topological features, persistence diagrams are commonly used. Persistence diagrams are visual representations used in the study of persistent homology to illustrate the birth and death of topological features (such as connected components, loops, and voids) as the scale of observation changes. It consists in a set of points $(b_{i},d_{i})\in\mathbb{R}^{2}$ (with $b_{i}\leq d_{i}$). Each point represents the birth ($b_{i}$) and death ($d_{i}$) of a topological feature in a filtered space. Instead of points, we sometimes draw intervals representing the birth and death times of homology classes. This method, also known as persistence barcodes [31], provides an alternative visualization of the same information contained in persistence diagrams, plotting this data in bars. Each bar in the barcode corresponds to a topological feature and is represented by an interval [$b_{i},d_{i}$). An illustrative example of a persistence diagram and a persistence barcode is in Fig. 2.
@@
Persistent entropy. In order to measure how much the construction of a filtered simplicial complex is ordered, an entropy measure, called persistent entropy, was defined in [7]. A precursor of this definition was given in [24] to measure how different bars of a persistence barcode are in length. In other words, persistent entropy is a measure of the complexity of a topological space based on its persistence diagram.
Given a filtered simplicial complex $\{K(t):t\in\mathbb{R}\}$, and the corresponding persistence diagram $D=\{(b_{i},d_{i}):i\in I\}$, where $b_{i}\leq d_{i}$ for all $i\in I$, and $I$ is the index set that identifies the pairs $(b_{i},d_{i})$ in $D$. As commented previously, each pair $(b_{i},d_{i})$ corresponds to a topological feature that “appears” at time $b_{i}$ (birth) and “disappears” at time $d_{i}$ (death) as the filtration progresses. The persistent entropy $H$ of a filtered simplicial complex $K$ is calculated as follows:
$H=-\sum_{i\in I}p_{i}\ln{p_{i}}$ (1)
@@
The maximum persistent entropy corresponds to the situation in which all the intervals in the barcode are of equal length. In that case, $H=\ln{n}$, where is the number of elements of $I$. In contrast, the value of the persistent entropy decreases as more intervals of different lengths are present. A potential point of confusion arises because persistent entropy does not measure the entropy of the distribution of bar lengths directly. Instead, it measures the entropy of the probability distribution induced by the relative contributions of each bar to the total barcode length. If all bars had the same length, their contributions would be equal, maximizing entropy. Conversely, when some bars dominate the total length, entropy decreases. Thus, persistent entropy should be interpreted as the entropy of the distribution of bar indices weighted by their respective lifetimes, rather than the entropy of the length distribution itself.
For example, for the persistence barcode of dimension 0 in Fig. 2(b), we have 5 bars with the following lengths $l_{i}$: $1,1,2,2,8$, obtaining a total length $L$ of $14$. Thus, the probabilities ($p_{i}$) in this case are: $\frac{1}{14}$, $\frac{1}{14}$, $\frac{2}{14}$, $\frac{2}{14}$, and $\frac{8}{14}$, resulting in a persistent entropy of 1.25. A different distribution of the same points producing lower or higher persistent entropy indicates that the points are more concentrated or dispersed, respectively.
@@ 3.2 Background on Confidence Regions
Adjustable Classifiers. Consider an input space $X\subseteq\mathbb{R}^{d}$ and an output space $Y=\{-1,+1\}$, where a binary adjustable classifier is defined as:
@@
Probabilistic Scaling. Probabilistic scaling (SP) constructs safety regions based on the order statistics of the calibration dataset $Z_{c}=\{(x_{i},y_{i})\}_{i=1}^{n^{c}}\subset X\times Y$. We define the generalized maximum $\text{max}_{(r)}(\Gamma)$ for a set $\Gamma=\{\gamma_{i}\}_{i=1}^{n}\in\mathbb{R}^{n}$ as the $r$-th largest value, ensuring no more than $r-1$ elements of $\Gamma$ exceed $\text{max}_{(r)}(\Gamma)$. Assuming a continuous and monotonically increasing $f_{\theta}(x,\rho)$ on $\rho$, probabilistic scaling yields a $\rho_{\epsilon}$ such that:
@@
Conformal Prediction. Conformal prediction (CP), as developed in [32], provides a post-hoc assessment of classification conformity. Using a score function $s:X\times Y\rightarrow\mathbb{R}$, such as $s(x,\hat{y})=-\hat{y}\bar{\rho}(x)$, that encodes the agreement between a sample $x$ and a candidate label $\hat{y}$, CP defines a prediction region $C_{\epsilon}(x)$:
@@ 3.3 Rule-Based Models Background
Global Rule Classifiers. In this kind of approach, the model learns a set of rules that represent the entire logic of the dataset, making it suitable for application to any data sample. The rules are “native” as they arise directly from the learning process without needing any intermediary steps.
Formally, a rule-based classifier trained on a dataset $T=\{(x_{j},y_{j})\}_{j=1}^{N}\in X\times Y$, where $x_{j}\in\mathbb{R}^{d}$ and $y_{j}\in\{-1,+1\}$, generates a ruleset =$\mathcal{R}=\{r_{k}\}_{k=1}^{M}$. Each rule $r_{k}$ has a premise, or antecedent, expressed as a conjunction of conditions:
@@
Local Rule Extraction via Anchors.Anchors is a model-agnostic local rule extraction technique that generates high-precision rules for explaining individual predictions of any black-box classifier. While locally faithful, these rules also hold in a neighborhood (or perturbation space) of the instance being explained. An anchor $A$ for an instance $x$ is a set of predicates that satisfies the precision threshold $\lambda_{\text{prec}}$ with a confidence level $1-\delta$:
@@
Rule Evaluation. For rule-based classifiers, the performance of each rule $r_{k}$ can be measured with two key metrics—coverage $C(r_{k})$ and error $E(r_{k})$—that allow us to evaluate the rule’s ability to generalize to unseen data. Before defining the metrics, it is worth underlining that a ‘positive’ instance, in this context, refers to any instance that satisfies the considered rule, regardless of the output class. Similarly, the term ‘negative’ is used to denote the case when points do not satisfy the rule.
@@ 3.4 Simulation of Social Robotics Navigation
Figure 3: Point cloud representation of robot positions in the “cross” scenario. In this setup, robots navigate between predefined waypoints(colored cylinders) along the x-axis and y-axis, creating intersecting flows at the center.
@@ 4 Topology-Driven Safety Methodology
      Normalize interval lengths: We compute the sum of all interval lengths and normalize each length by this sum, obtaining a probability distribution of interval lengths.
@@
      Calculate entropy: We compute the entropy of the probability distribution, using the Shannon entropy formula Eq. (1). See Fig. 5 for an example of a point cloud corresponding to the position of the robots in a specific time step, the corresponding persistence barcode of dimension 0, and the corresponding persistent entropy.
@@
      Generate time series and statistical parameters. Once the above process is done for each time step of the simulation, we obtain a persistent-entropy time series for the whole simulation. Finally, we compute four key statistical parameters for this time series: mean, median, standard deviation and interquartile range (IQR). See Fig. 6 for an example of persistent entropy time series over a 500-step simulation.
@@
    • Input $x$: A feature vector extracted from the topological data of the simulation, defined as: $x=(meanEntropy,medianEntropy,stdEntropy,iqrEntropy)$. These features are derived from the 6-step methodology explained above.
@@
      $y_{i}=\begin{cases}+1&\text{if the number of collisions during the simulation $i$ equals $0$,}\\ -1&\text{if the number of collisions during the simulation $i$ is greater than$0$}.\end{cases}.$
@@
Note that a specific time step in a simulation with a higher persistent entropy indicates a greater dispersion among the robots in relation how they are distributed around the space, without clusters of robots in certain areas and without empty areas or areas with few robots, being less likely to have empty areas and to form cluster of robots very grouped and another ones very far away, while a lower entropy indicates that less dispersion between the robots, being more likely to occupy specific regions of space and form clusters. See Fig. 7 for an illustrative example, where we can compare the robot point cloud of two different time steps, the point cloud on the left with lower entropy than the one on the right (where the points are more dispersed).
@@ 5 Experiments
In this section, we present the experiments and results we carried out by putting together the simulator, the safety regions and the illustrated topological techniques. After data collection through Navground simulator, we study the problem of collision avoidance, comparing our methodology with the one that uses behavior parameters ($sigma,\eta\text{ and }\tau$, which are defined in Tab. 1 for the construction of safety regions [17] (instead of the topological features), then studying the problem of deadlock avoidance, and finally combining them for safe and efficient (free of collisions and deadlocks) navigations.
@@ 5.1 Data Collection
For each simulation run, we apply the methodology explained previously (Section 4), recording the following values
$\textbf{x}=(meanEntropy,medianEntropy,stdEntropy,iqrEntropy),$
@@ 5.3 Results
First, we will compare training a rule-based model with the behavior parameters (as is done in [17]) against training it with the topological features extracted with this methodology for collision avoidance. For that, we will use Skope Rules. Skope Rules [^4] https://github.com/scikit-learn-contrib/skope-rules is a rule-based ML algorithm, that learns interpretable and diversified rules for “scoping” a target class of interest, i.e. detecting samples from this class with high precision. In practical applications, skope-rules model has been applied to several safety-critical classification tasks, as well as anomaly detection problems or cluster description.
@@
SkopeRules-Behavior Parameters 8 83.6 81.9 92.0 27.5 8.0 72.5
SkopeRules-Topological Features $20$ $84.6$ $84.6$ $87.1$ $18.7$ $12.9$ $81.3$
Table 2 highlights the superior performance of SkopeRules using topological features compared to behavior parameters. The topological model achieves higher accuracy (84.6% vs. 83.6%), F1-score (84.6% vs. 81.9%), and true negative rate (81.3% vs. 72.5%) while significantly reducing the false positive rate (18.7% vs. 27.5%). Although its true positive rate (87.1% vs. 92.0%) is slightly lower, and the false negative rate (12.9% vs. 8.0%) is higher, these trade-offs are outweighed by its improved balance across metrics. Remember that the positive class is the non-collision class (+1).
The use of 20 rules, compared to 8 for behavior parameters, suggests the topological model captures more complex relationships, enhancing its robustness. Models with fewer rules have the advantage of being more interpretable, but the richest ones may generate more fine-grained rules with better discriminative ability. Overall, topological features provide a more discriminative representation, making this model preferable for scenarios prioritizing accuracy and the minimization of false positives.
Also, we will compare the safety regions generated using these two types of input features. For that, the techniques detailed in Section 3.2 to derive the safety regions using Conformal Prediction and Probabilistic Scaling were applied by adopting an SVM as the classifier. Specifically, we considered a Gaussian kernel SVM, with a sigma set to 0.5, with regularization parameter set to 0.3, and a weighting of 0.5. With this base model, prior to any error control, we achieve the following performance: 78.2% accuracy, 80.6% F1 Score, 82.8% TPR, 72.7% TNR, 27.3% FPR, 17.2% FNR using behavior parameters as input and 85% accuracy, 88.1% F1 Score, 91.6% TPR, 75% TNR, 25% FPR, 8.4% FNR using topological features as input.
@@
Using the SkopeRules method and topological features to classify simulations with deadlock cases versus efficient simulations without them, we derive a total of 18 rules, achieving highly positive performance: 89% accuracy, 88.5% F1-score, 93% TPR, and 84% TNR. This demonstrates that these features achieve even better performance for deadlock avoidance than for collision avoidance when using SkopeRules.
Next, we generate the safety regions using these features in the same manner as before. We considered a Gaussian kernel SVM, with a sigma set to 0.5, a regularization parameter set to 0.3, and a weighting of 0.5. With this base model, prior to any error control, we achieve the following performance: 86% accuracy, 82.3% F1 score, 74.7% TPR, 94.7% TNR. Using PS method we achieve the following performance: $\rho_{\epsilon}$ equal to -0.12, 87.2% accuracy, 84.6% F1 score, 80.1% TPR, 92.7% TNR; and using the CP method we achieve the following performance: $\rho_{\epsilon}$ equal to 0.17, 83.6% accuracy, 77.9% F1 score, 66% TPR, 97.3% TNR.
@@
- 91% coverage and 21% error for ground truth labels, and 100% coverage and 22% for method labels for PS method.
- 91% coverage and 21% error for ground truth labels, and 100% coverage and 39% for method labels for PS method.
We can see how with the PS method we obtain a better performance in terms of error, and with the CS method a better performance in terms of covering (as for collision avoidance using topological features), although the rules are exactly the same, obviously obtaining the same performance for the real labels. Considering the non-trivial nature of the problem and of approximating a complex SVM shape via hyper-rectangular shapes (rules) while keeping the error bound as low as possible, we can consider our results as a promising compromise between safety and transparency.

### v2

@@ 2 Related work
Moreover, persistent entropy has demonstrated its effectiveness in a variety of other applications such as characterizing idiotypic immune networks [25], analyzing similarities in piecewise linear functions and time series [7], and separating topological features from noise in Vietoris-Rips complexes [26]. These successful applications underscore its versatility and potential to provide robust insights across diverse domains. Due to its simplicity (providing a single scalar value at each time step making it computationally efficient while still capturing essential topological features), interpretability, stability [27], and proven success in other applications, as previously mentioned, make it an ideal choice for our work.
To the best of our knowledge, there are no previous works that utilize TDA-based methods specifically for defining safety regions in robot navigation. We have reviewed the state of the art and found that TDA has indeed been explored in various aspects of robot navigation, particularly for characterizing interactions and spatial structures (e.g., [5, 6, 21, 22, 23]). However, none of these works explicitly address the definition of safety regions using TDA.
@@ 3.1 Topology Background
3.1.1 Topology and Topological Space.
A topological space is a powerful mathematical concept for describing the connectivity of a space. Informally, a topological space is a set of points, each of them equipped with the notion of neighbourhood.
@@
3.1.2 Simplicial Complex.
An abstract simplicial complex $\mathcal{K}$ is given by:
  • A set $V$ of 0-simplices (also called vertices);
@@
A simplicial complex $K$ is a geometric realization of an abstract simplicial complex $\mathcal{K}$ [30]. It is constructed as a nested family of simplices, where each simplex is a generalization of geometric shapes of various dimensions. Specifically, a 0-simplex corresponds to a point in a given Euclidean space $\mathbb{R}^{n}$, a $1$-simplex represents an edge, a $2$-simplex is a filled triangle, a $3$-simplex is a filled tetrahedron, and so on. This hierarchical structure allows the simplicial complex to model higher-dimensional relationships between points in a mathematically rigorous way.
Homology groups are algebraic structures that describe features of a topological space $\mathcal{C}$. The $k$-th Betti number represents the rank of the $k$-dimensional homology group. Informally, for a fixed $k$, the $k$-th Betti number $\beta_{k}$ counts the number of $k$-dimensional holes characterizing $\mathcal{C}$: $\beta_{0}$ is the number of connected components, $\beta_{1}$ counts the number of holes in $2D$ or tunnels in $3D$ 111dDdD refers to the $d$ - dimensional space $\mathbb{R}^{d}$, $\beta_{2}$ can be thought of as the number of voids in geometric solids. In this work, we use simplicial homology, which consists of computing homology groups of simplicial complexes.
@@
3.1.3 Persistent Homology.
Topological data analysis, particularly persistent homology, has emerged as a valuable approach for studying the geometry and connectivity of datasets that evolve over time or space [28]. The primary objective of employing topological data analysis tools, like persistent homology and persistent entropy, in the context of fleet behavior modeling is to achieve a more nuanced understanding of system dynamics.
@@
Persistent homology describes how the homology of $K$ changes along a filtration $\{K(t)\mid t\in\mathbb{R}\}$. A $k$-dimensional Betti interval, with endpoints $t_{\text{start}},t_{\text{end}})$, corresponds to a $k$-dimensional hole that appears at filtration time $t_{\text{start}}$ and remains until time $t_{\text{end}}$.
To visualize the persistence of topological features, persistence diagrams are commonly used. Persistence diagrams are visual representations used in the study of persistent homology to illustrate the birth and death of topological features (such as connected components, loops, and voids) as the scale of observation changes. It consists in a set of points $(b_{i},d_{i})\in\mathbb{R}^{2}$ (with $b_{i}\leq d_{i}$). Each point represents the birth ($b_{i}$) and death ($d_{i}$) of a topological feature in a filtered space. Instead of points, we sometimes draw intervals representing the birth and death times of homology classes. This method, also known as persistence barcodes [32], provides an alternative visualization of the same information contained in persistence diagrams, plotting this data in bars. Each bar in the barcode corresponds to a topological feature and is represented by an interval $b_{i},d_{i}$). An illustrative example of a persistence diagram and a persistence barcode is in Fig. 2.
@@
3.1.4 Persistent Entropy.
In order to measure how much the construction of a filtered simplicial complex is ordered, an entropy measure, called persistent entropy, was defined in [7]. A precursor of this definition was given in [24] to measure how different bars of a persistence barcode are in length. In other words, persistent entropy is a measure of the complexity of a topological space based on its persistence diagram.
Given a filtered simplicial complex $\{K(t):t\in\mathbb{R}\}$, and the corresponding persistence diagram $D=\{(b_{i},d_{i}):i\in I\}$, where $b_{i}\leq d_{i}$ for all $i\in I$, and $I$ is the index set that identifies the pairs $(b_{i},d_{i})$ in $D$. As commented previously, each pair $(b_{i},d_{i})$ corresponds to a topological feature that “appears” at time $b_{i}$ (birth) and “disappears” at time $d_{i}$ (death) as the filtration progresses. The persistent entropy $\operatorname{PE}$ of a filtered simplicial complex $K$ is calculated as follows:
$\operatorname{PE}=-\sum_{i\in I}p_{i}\ln{p_{i}}$ (1)
@@
The maximum persistent entropy corresponds to the situation in which all the intervals in the barcode are of equal length. In that case, $\operatorname{PE}=\ln{n}$, where $n$ is the number of elements of $I$ (i.e, the number of bars in the persistence diagram $D$). In contrast, the value of the persistent entropy decreases as more intervals of different lengths are present. A potential point of confusion arises because persistent entropy does not measure the entropy of the distribution of bar lengths directly. Instead, it measures the entropy of the probability distribution induced by the relative contributions of each bar to the total barcode length. If all bars had the same length, their contributions would be equal, maximizing entropy. Conversely, when some bars dominate the total length, entropy decreases. Thus, persistent entropy should be interpreted as the entropy of the distribution of bar indices weighted by their respective lifetimes, rather than the entropy of the length distribution itself.
For example, for the persistence barcode of dimension 0 in Fig. 2(b), we have 5 bars (red bars correspond to 0 dimension topological features) with the following lengths $l_{i}$: $1,1,2,2,8$, obtaining a total length $L$ of $14$. Thus, the probabilities ($p_{i}$) in this case are: $\frac{1}{14}$, $\frac{1}{14}$, $\frac{2}{14}$, $\frac{2}{14}$, and $\frac{8}{14}$, resulting in a persistent entropy of 1.25.
The stability insights derived from persistence diagrams [33] and persistent entropy [27, 7] offer a solid foundation for measuring the stability of topological features in response to minor changes. In particular, they guarantee that small perturbations in the input lead to short bars in the persistence barcode and result in minimal, bounded shifts in the persistent entropy. This property is especially useful in contexts like robot navigation, where slight variations in the robot’s position—such as those occurring between two time instances with minimal displacement—can otherwise lead to significant uncertainties in the system’s analysis.
@@ 3.2 Background on Confidence Regions
3.2.1 Adjustable Classifiers.
Consider an input space $X\subseteq\mathbb{R}^{d}$ and an output space $Y=\{-1,+1\}$, where a binary adjustable classifier is defined as:
@@
3.2.2 Probabilistic Scaling.
Probabilistic scaling (SP) constructs safety regions based on the order statistics of the calibration dataset $Z_{c}=\{(x_{i},y_{i})\}_{i=1}^{n^{c}}\subset X\times Y$. We define the generalized maximum $\text{max}_{(r)}(\Gamma)$ for a set $\Gamma=\{\gamma_{i}\}_{i=1}^{n}\in\mathbb{R}^{n}$ as the $r$-th largest value, ensuring no more than r1r-1 elements of $\Gamma$ exceed $\text{max}_{(r)}(\Gamma)$. Assuming a continuous and monotonically increasing $f_{\theta}(x,\rho)$ on $\rho$, probabilistic scaling yields a $\rho_{\epsilon}$ such that:
@@
3.2.3 Conformal Prediction.
Conformal prediction (CP), as developed in [34], provides a post-hoc assessment of classification conformity. Using a score function $s:X\times Y\rightarrow\mathbb{R}$, such as $s(x,\hat{y})=-\hat{y}\bar{\rho}(x)$, that encodes the agreement between a sample $x$ and a candidate label $\hat{y}$, CP defines a prediction region $C_{\epsilon}(x)$:
@@ 3.3 Rule-Based Models Background
3.3.1 Global Rule-Based Classifiers.
In this kind of approach, the model learns a set of rules that represent the entire logic of the dataset, making it suitable for application to any data sample. The rules are “native” as they arise directly from the learning process without needing any intermediary steps.
Formally, a rule-based classifier trained on a dataset $T=\{(x_{j},y_{j})\}_{j=1}^{N}\in X\times Y$, where $x_{j}\in\mathbb{R}^{d}$ and $y_{j}\in\{-1,+1\}$, generates a ruleset $\mathcal{R}=\{r_{k}\}_{k=1}^{M}$. Each rule $r_{k}$ has a premise, or antecedent, expressed as a conjunction of conditions:
@@
3.3.2 Local Rule Extraction via Anchors.
Anchors is a model-agnostic local rule extraction technique that generates high-precision rules for explaining individual predictions of any black-box classifier. While locally faithful, these rules also hold in a neighborhood (or perturbation space) of the instance being explained. An anchor $A$ for an instance $x$ is a set of predicates that satisfies the precision threshold $\lambda_{\text{prec}}$ with a confidence level $1-\delta$:
@@
3.3.3 Rule Evaluation.
For rule-based classifiers, the performance of each rule $r_{k}$ can be measured with two key metrics—coverage $C(r_{k})$ and error $E(r_{k})$—that allow us to evaluate the rule’s ability to generalize to unseen data. Before defining the metrics, it is worth underlining that a ‘positive’ instance, in this context, refers to any instance that satisfies the considered rule, regardless of the output class. Similarly, the term ‘negative’ is used to denote the case when points do not satisfy the rule.
@@ 3.4 Simulation of Social Robotics Navigation
Figure 3: In the crossings scenario, simulated robots navigate between predefined waypoints (colored cylinders) along the x-axis and y-axis, creating intersecting flows at the center.
@@ 4 Topology-Driven Safety Methodology
      Normalize interval lengths: We compute the sum of all interval lengths and normalize each length by this sum, obtaining a probability distribution of interval lengths of the topological features.
@@
      Calculate persistent entropy: We compute the entropy of the topological features probability distribution, using the Shannon entropy formula Eq. (1). See Fig. 5 for an example of a point cloud corresponding to the position of the robots in a specific time step, the corresponding persistence barcode of dimension 0, and the corresponding persistent entropy.
@@
      Generate persistent entropy time series and compute statistical parameters. Once the above process is done for each time step of the simulation, we obtain a persistent entropy time series for the whole simulation. Finally, we compute four key statistical parameters for this time series: mean, median, standard deviation and interquartile range (IQR). See Fig. 6 for an example of persistent entropy time series over a 500-step simulation.
@@
    • Input $x$: A feature vector extracted from the topological data of the simulation, defined as: x=(meanEntropy,medianEntropy,stdEntropy,iqrEntropy)x=(meanEntropy,medianEntropy,stdEntropy,iqrEntropy). These features are derived from the 6-step methodology explained above.
@@
[…]
@@
Note that a specific time step in a simulation with a higher persistent entropy indicates a greater dispersion among the robots in relation how they are distributed around the space, without clusters of robots in certain areas and without empty areas or areas with few robots, being less likely to have empty areas and to form cluster of robots very grouped and another ones very far away, while a lower entropy indicates that less dispersion between the robots, being more likely to occupy specific regions of space and form clusters. See Fig. 7 for an illustrative example where we can compare the robot point cloud from two different time steps. The point cloud on the left has lower entropy than the one on the right (where the points are more dispersed).
@@ 5 Experiments
In this section, we present the experiments and results we carried out by putting together the simulator, the safety regions and the illustrated topological techniques. After data collection through Navground simulator, we study the problem of collision avoidance, comparing our methodology with the one that uses behavior parameters ($\sigma,\eta\text{ and }\tau$, which are defined in Tab. 1 for the construction of safety regions [17] (instead of the topological features), then studying the problem of deadlock avoidance, and finally combining them for safe and efficient (free of collisions and deadlocks) navigations.
@@ 5.1 Data Collection
For each simulation run, we apply the methodology explained previously (Section 4), recording the following values:
$\textbf{x}=(meanEntropy,medianEntropy,stdEntropy,iqrEntropy).$
According to the computational cost of calculating persistent entropy, for instance, the calculation time for persistent entropy for a point cloud at a given time $i$ is less than 10 ms on a single core of a modern CPU (approximately 0.003 seconds), and the total time for computing the entire persistent entropy time series for a simulation with the mentioned characteristics is approximately 0.2 seconds. These times suggest that the method is efficient enough for real-time performance in practical applications.
@@ 5.3 Results
First, we will compare training a rule-based model with the behavior parameters (as is done in [17]) against by training it with the topological features that have been extracted using the methodology proposed in this paper for the avoidance of collisions. For that, we will use Skope Rules. Skope Rules [^4] https://github.com/scikit-learn-contrib/skope-rules is a rule-based ML algorithm, that learns interpretable and diversified rules for “scoping” a target class of interest, i.e. detecting samples from this class with high precision. In practical applications, skope-rules model has been applied to several safety-critical classification tasks, as well as anomaly detection problems or cluster description.
@@
SkopeRules-Behavior Parameters 99 $78.7$ $75.1$ $70.8$ $15$ $29.2$ $85$
SkopeRules-Topological Features 2020 $84.6$ $84.6$ $87.1$ $18.7$ $12.9$ $81.3$
Table 2 highlights the superior performance of SkopeRules using topological features compared to behavior parameters. For a proper interpretation of the results, we remark that, based on the labeling criteria defined in Sec. 5.1, the positive class (i.e., $y=+1$) refers to our target situation (that is, either absence of collisions, deadlocks, or absence of both). As a result, true positives here denote their correct classification, with false positives reflecting how many hazardous cases (presence of collisions, deadlocks, or both, labeled as y=1y=-1) are missed. In contrast, true negatives are associated to the dangerous situations being correctly classified, with false negatives representing false alarm cases.
The model using topological features as input features achieves higher accuracy (84.6% vs. 78.6%), F1-score (84.6% vs. 75.1%), and true positive rate (87.1% vs. 70.8%) while significantly reducing the false negative rate (12.9% vs. 29.2%). Although its true negative rate (81.3% vs. 85%) is slightly lower, and the false positive rate (18.7% vs. 15%) is higher, these trade-offs are outweighed by its improved balance across metrics. Remember that the positive class is the non-collision class (+1).
The use of 20 rules, compared to 9 for behavior parameters, suggests that the model that using topological features as input features, captures more complex relationships, enhancing its robustness. Models with fewer rules have the advantage of being more interpretable, but the richest ones may generate more fine-grained rules with better discriminative ability. Overall, topological features provide a more discriminative representation, making the model using them as input features preferable for scenarios where accuracy and minimising false positives are prioritised.
Also, we will compare the safety regions generated using these two types of input features. For that, the techniques detailed in Section 3.2 to derive the safety regions using Conformal Prediction and Probabilistic Scaling were applied by adopting an SVM as the classifier. Specifically, we considered a Gaussian kernel SVM, with a sigma set to 0.5, with regularization parameter set to 0.3, and a weighting of 0.5. These hyperparameters were selected manually to ensure a fair and controlled comparison across the different types of input features, rather than focusing on obtaining the best possible SVM performance. With this base model, prior to any error control, we achieve the following performance: 78.2% accuracy, 80.6% F1 Score, 82.8% TPR, 72.7% TNR, 27.3% FPR, 17.2% FNR using behavior parameters as input and 85% accuracy, 88.1% F1 Score, 91.6% TPR, 75% TNR, 25% FPR, 8.4% FNR using topological features as input.
@@
Using the SkopeRules method and topological features to classify simulations with deadlock cases versus efficient simulations without them, we derive a total of 18 rules, achieving highly positive performance: 88.6% accuracy, 89% F1-score, 93% TPR, and 84% TNR. This demonstrates that these features achieve even better performance for deadlock avoidance than for collision avoidance when using SkopeRules.
Next, we generate the safety regions using these features in the same manner as before. We again considered a Gaussian kernel SVM, with a sigma set to 0.5, a regularization parameter set to 0.3, and a weighting of 0.5. With this base model, prior to any error control, we achieve the following performance: 86% accuracy, 82.3% F1 score, 74.7% TPR, 94.7% TNR. Using PS method we achieve the following performance: $\rho_{\epsilon}$ equal to -0.12, 87.2% accuracy, 84.6% F1 score, 80.1% TPR, 92.7% TNR; and using the CP method we achieve the following performance: $\rho_{\epsilon}$ equal to 0.17, 83.6% accuracy, 77.9% F1 score, 66% TPR, 97.3% TNR.
@@
- 91% coverage and 21% error for ground truth labels, and 100% coverage and 22.5% for method labels for PS method.
- 91% coverage and 21% error for ground truth labels, and 100% coverage and 40% for method labels for CP method.
We can see how with the PS method we obtain a better performance in terms of error, although the rules are exactly the same, obviously obtaining the same performance for the real labels. Considering the non-trivial nature of the problem and of approximating a complex SVM shape via hyper-rectangular shapes (rules) while keeping the error bound as low as possible, we can consider our results as a promising compromise between safety and transparency.